Marsha has money in two savings accounts. One rate is 8% and the other is 10%. If she has $200 more in the 10% account, how much is invested at 8% if the total interest is $101?
Need a little help with this problem, Thank you!

Answer by checkley71(8403)

.08X+.10(X+200)=101
.08X+.10X+20=101
.18X=101-20
.18X=81
X=81/.18
X=450
PROOF
.08*450+.10(450+200)=101
36+.10(650)=101
36+65=101
101=101